2. A researcher claims that the average salary of a private school teacher is greater than P20,000 with a standard deviation of P5,000. A sample of 20 teachers has a mean salary of P25,000. Test the claim of the researcher. At 0.05 level of significance, test the claim of the researcher.



Answer the following:



Parameter:


Claim:


Claim ( in symbol ):


Ho: Ho:


Ha: Ha:


What is the significance level or a?


Is it two-tailed or one-tailed test?

1
Expert's answer
2022-04-29T13:29:52-0400

Parameter: mean

Claim: the average salary of a private school teacher is greater than P20,000 

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le20000"

"H_a:\mu>20000"

One-tailed test.

Significance level is "\alpha = 0.05."

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a right-tailed test is "z_c = 1.6449."

The rejection region for this right-tailed test is "R = \{z: z > 1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{25000-20000}{5000/\sqrt{20}}\approx4.4721"

Since it is observed that "z = 4.4721>1.6449= z_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p=P(Z>4.4721)=0," and since "p=0<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than 20000, at the "\alpha = 0.05" significance level.